# Heterogeneous Treatment Effects and

### v3

@@ 3.1 Causal Effects
We use $ADE$ and $AIE$ to denote average direct effect and average indirect effect. Throughout the paper, we assume this expectation is well-defined. Our framework proceeds by linking the indirect effect associated with a mechanism to heterogeneous treatment effects with respect to a covariate. To this end, define conditional average treatment effects as follows:
Definition 1 (Conditional Average Treatment Effect).
Consider pre-treatment covariate $X_{k}$. Given that $z\neq z^{\prime}\in Z$, the conditional average treatment effect (CATE) with respect to $X_{k}=x$ is:
$CATE(X_{k}=x)=E_{X_{\neg k}}[Y(Z=z,M;X)-Y(Z=z^{\prime},M;X)|X_{k}=x]$
There are many ways to define and measure treatment effect heterogeneity. In this article, we consider HTEs with respect to pre-treatment moderators. This adheres to the common practice of using HTEs to detect mechanisms that we documented in Table 1.
Definition 2 (Heterogeneous treatment effects).
HTEs exist with respect to pre-treatment covariate $X_{k}$ if $CATE(X_{k}=x)\neq CATE(X_{k}=x^{\prime})$ for some $x\neq x^{\prime}\in X_{k}$. [^11] More precisely, the probability measure for the set that contains such $x$ and $x^{\prime}$ is non-zero.
With this framework, it is now possible to express the research question more precisely. First, recall our research question: “under what conditions do HTEs with respect to pre-treatment covariates provide evidence of mechanism activation?” By mechanism activation, we mean that the indirect effect of a mechanism is non-zero for some unit. The use of heterogeneous treatment effects registers the expectation that the mechanism need not be active for all units in the population. Our question then can be stated more precisely: “Under what conditions are HTEs with respect to a covariate $X_{k}$ sufficient to show that there exists some unit for which $IE(z,z^{\prime};X_{k})\neq 0$?”
3.2 Mechanisms and Outcomes
In standard discussions of research design, researchers generally do not make methodological distinctions between different types of outcomes. We introduce a distinction between outcomes that are directly affected and indirectly affected by a mechanism. In the theoretical model of our motivating example and many related theories in this literature, the mechanism—voter learning—acts directly on voter utility. The mechanism indirectly affects vote choice through its (assumed) effect on utility. Definition 3 shows that the distinction between directly- and indirectly-affected outcomes rests on fundamentally on the theorized causal sequence of different outcomes of interest. The sequencing of multiple outcomes has widespread implications for efforts to identify and measure causal effects, but has received limited attention in the existing literature (for one exception, see Slough, 2023).
@@ 3.2 Mechanisms and Outcomes
Given treatment $Z$, let variable $\widetilde{Y}(Z)$ be a (potential) outcome. If
@@
    There exists another (potential) outcome $Y(Z)$ that causally precedes $\widetilde{Y}$; and

### v4

@@ 2.2 Causal Mediation and Indirect Effects
We use $ADE^{Y}$ and $AIE^{Y}_{j}$ to denote average direct effect and average indirect effect of mechanism $j$ for outcome $Y$, respectively. Throughout the paper, we assume the expectation in (10) is well-defined.
2.3 HTEs and the Identification of Indirect Effects
How do HTEs—differences in CATEs—relate to the indirect effects that are estimated within the mediatiation framework? To show this relationship, consider a decomposition of a CATE into a conditional ADE and two conditional AIEs following (11) [^10] Conditional $ADE^{Y}(z,z^{\prime};X_{k}=x)$ and $AIE_{j}^{Y}(z,z^{\prime};X_{k}=x)$ evaluate expectations over all $X$ but fixing $X_{k}$ at $x$.:
$CATE^{Y}(X_{k}=x)=ADE^{Y}(z,z^{\prime};X_{k}=x)+\sum_{j=1}^{2}AIE^{Y}_{j}(z,z^{\prime};X_{k}=x)$
One can then express the difference in CATEs, at two distinct levels of $X_{k}$ as:
$\begin{aligned} CATE^{Y}(X_{k}=x)-CATE^{Y}(X_{k}=x^{\prime})=&[ADE^{Y}(z,z^{\prime};X_{k}=x)-ADE^{Y}(z,z^{\prime};X_{k}=x^{\prime})]\\ &+\sum_{j=1}^{2}[AIE^{Y}_{j}(z,z^{\prime};X_{k}=x)-AIE^{Y}_{j}(z,z^{\prime};X_{k}=x^{\prime})]\end{aligned}$ (12)
From (12), it is clear that HTEs could arise from differences in direct and/or indirect effects. Suppose that we were interested in evaluting the activation of mechanism 1 with respect to outcome $Y$. Expression (12) shows that we cannot automatically attribute observed heterogeneity to mechanism 1. Instead, to link a difference in CATEs (HTEs) to a difference in indirect effects, we need to assume that the conditional direct effect $ADE^{Y}(z,z^{\prime};X_{k}=x)$ and any conditional indirect effect(s) of the other mechanism(s) $AIE^{Y}_{2}(z,z^{\prime};X_{k}=x)$ do not vary at different levels of $X_{k}$, as stated in Assumptions 1-2. Assumption 1 formalizes an exclusion assumption posited by Imai et al. (2011), who focus on the case of a single mechanism. [^11] Specifically, Imai et al. (2010a: p. 784) write “if the size of the ADE does not depend on the pretreatment covariate … a statistically significant interaction term implies that the [average causal mediation effect] is larger for one group … than for another group.”
Assumption 1 (Exclusion I).
Given $z,z^{\prime}\in Z$ and $x,x^{\prime}\in X_{k}$, $X_{k}$ is excluded from the direct effect such that $ADE^{Y}(z,z^{\prime};X_{k}=x)=ADE^{Y}(z,z^{\prime};X_{k}=x^{\prime})$.
Assumption 2 (Exclusion II).
Given $z,z^{\prime}\in Z$ and $x,x^{\prime}\in X_{k}$, $X_{k}$ is excluded from the indirect effect of the other mechanism 2: $AIE^{Y}_{2}(z,z^{\prime};X_{k}=x)=AIE^{Y}_{2}(z,z^{\prime};X_{k}=x^{\prime})$.
Assumptions 1-2 constrain the relationship between a moderator, $X_{k}$, other mechanisms (e.g., $M_{2}$) and any direct effect of treatment. [^12] Other methods for mechanism detection invoke distinct exclusion assumptions. For example, the scaling stage of implicit mediation uses instrumental variables analysis to estimate causal mediation effects (Bullock and Green, 2021). In contrast to our analysis of HTEs, the exclusion assumptions underpinning implicit mediation restrict the mediators that might be activated by a treatment. See also Fu (2024) for a related discussion on mediation analysis with HTEs. Figure 1 illustrates these assumptions graphically. While this figure resembles a directed acyclic graph (DAG), we depart from conventional presentation of DAGs as vertices (nodes) and edges (arrows) because there are edges that point to other edges (rather than vertices). We make this departure because our assumptions impose greater structure on the possible causal moderation (or lack thereof) than is assumed in traditional DAGs. This departure is not new. Nilsson et al. (2021) notes that there is no standardized representation of causal moderation in DAGs, so our graphs are informed by an existing proposal for the representation of these effects by Weinberg (2007). This notation allows us to accurately convey the structure of causal moderation.
Once Assumptions 1 and 2 are invoked, it is straightforward to see that difference in CATEs reduces to differences in $AIE^{Y}_{1}$ at different levels of $X_{k}$, which we state formally in Proposition 1. Assumptions 1-2 are generically necessary because it is possible that $ADE^{Y}(x)-ADE^{Y}(x^{\prime})\neq 0$ and $AIE^{Y}_{2}(x)-AIE^{Y}_{2}(x^{\prime})\neq 0$ exactly offset each other. But under generic parameter values, the probability of this knife-edge event is zero.
[…]
Figure 1: Assumption 1 rules out the blue dashed path. Assumption 2 rules out both of the red dot-dashed paths. All black solid paths are permissible under Assumptions 1 and 2.
Proposition 1.
Assumptions 1-2 are sufficient and generically necessary for $CATE^{Y}(X_{k}=x)-CATE^{Y}(X_{k}=x^{\prime})=AIE^{Y}_{1}(z,z^{\prime};X_{k}=x)-AIE^{Y}_{1}(z,z^{\prime};X_{k}=x^{\prime})$.
Proposition 1 clarifies that a difference in CATEs does not identify either conditional AIE of mechanism 1 ($AIE^{Y}_{1}(z,z^{\prime};X_{k}=x)$ or $AIE^{Y}_{1}(z,z^{\prime};X_{k}=x^{\prime})$) in the absence of further assumptions. Rather, the difference in CATEs identifies a difference in conditional AIE’s. Thus, identification of this difference is not sufficient to identify indirect effects, as is the goal in (standard) mediation analysis. However, it is straightforward to see that if the difference in AIEs is not equal to zero, there must exist some unit for whom the AIE is not equal to zero. A non-zero difference in AIEs is therefore a sufficient condition for the activation of the relevant mechanism for at least one unit. This identification result motivates a more precise version of our research question: “Under what conditions are HTEs with respect to a covariate $X_{k}$ sufficient to show that there exists some unit for which $IE^{Y}_{1}(z,z^{\prime};X_{k})\neq 0$?”
To understand our later results, it is useful to see how Assumptions 1 and 2 could be violated. The first and most obvious violation would be that a covariate $X_{k}$ moderates multiple mechanisms (or one mechanism and the direct effect). This is clear from Figure 1. A second and less obvious violation is that the additive separability of the (indirect) effects of mechanisms breaks down. This would mean that if some $X_{k}$ moderates the effect of mechanism $1$, then it must also moderate the effect of any other mechanism that “interacts with” or whose (indirect) effect depends on the effect of mechanism $1$. In Figure 1, this would occur if, for example, there was an interaction between the effects of $M_{1}$ and $M_{2}$ on the outcome $Y$.
It is important to note that mediation analysis does not invoke Assumption 1 or 2, and instead invokes an assumption of sequential ignorability (Imai et al., 2010b). There is no logical ordering of the two types of assumptions: the exclusion assumptions do not imply sequential ignorability, nor does sequential ignorability imply the exclusion assumptions. This means that HTEs cannot said to be a more or less agnostic test of mechanism activation than mediation. In some applications, one set of assumptions may be more plausible or defensible than the other, but we cannot make a general claim about the strength of these distinct sets of assumptions. We provide a broader discussion comparing the use of HTEs to mediation analysis in E.
2.4 Connecting Mechanisms to Measured Variables
While our identification results show that HTEs can provide information about mechanism activation for some unit(s), the framework does not yet elucidate the relationship between a mechanism and measured variables. To understand why it is critical to develop this relationship beyond our identification results, note that a randomly-generated covariate that is independent of all variables in a research design (or system) satisfies Assumptions 1 and 2 by construction. Here, we would not expect to observe HTEs at different levels of the randomly-generated variable. But this lack of heterogeneity should not be informative about the substantive mechanism(s) at play.
We view a mechanism as an underlying process that responds to some activation and produces a given set of outputs. In the context of our running example, the voter distaste mechanism is activated by the observation of corruption information about the incumbent. It produces a number of outputs—a voter’s assessment of the incumbent and their voting decision—among other (unmodeled) possibilities. None of these objects—a voter’s information, their utility, or their voting decision—needs be inherently quantitative (though they could be). As social scientists, we choose how to measure and operationalize each of these objects: we normalize the voter information treatment to a binary (0/1) scale for convenient estimation of treatment effects; we choose some type of Likert scale to measure voter assessments/utility; and we choose self-reported vote choice or aggregated voting results to measure vote decisions. These operationalizations facilitate our ability to measure or quantify the effect of a mechanism on an outcome using, for example, a difference in CATEs. This implies that our ability to observe a mechanism’s effect depends fundamentally on how we choose to measure it (Slough and Tyson, 2024). Our concern here is therefore the link between a substantive mechanism and the measured variables in our framework.
2.4.1 Outcome variables and mechanisms
First, consider the relationship between measured outcomes and the outputs of a mechanism. A given mechanism produces multiple possible outputs; a researcher chooses to operationalize and measure some subset of those outputs as outcome variables. But not all outcome variables relate to the underlying mechanism in the same way. For example, in our motivating example the mechanism—voter distaste for observed corruption—affects a voter’s utility from the incumbent ($y_{1}$). The second outcome, vote choice, is a deterministic but non-linear function of utility given by the function $y_{2}(c)=\mathbb{I}[y_{1}(c)\geq 0]$. We refer to $y_{2}(c)$ as a transformed (potential) outcome relative to $y_{1}$.
@@ 2.4.1 Outcome variables and mechanisms
Given a (potential) outcome $Y(Z,M;X)$. Let $\widetilde{Y}(Z,M;X)=h(Y)$, where $h(\cdot)$ is a non-linear function. Then we call $\widetilde{Y}(Z,M;X)$ a transformed (potential) outcome relative to $Y(Z,M;X)$.
Formally, for a function $h(\cdot)$, we define the increment $\delta(t;d)=h(t+d)-h(t)$. We say that the function $h$ is non-linear if there exists nondegenerate $t,t^{\prime}$, and $d\neq 0$, such that the increment is different, $\delta(t;d)\neq\delta(t^{\prime};d)$. [^13] We assume that $h(t+d)$ is well defined. This is a general definition that does not assume differentiability of the function $h(\cdot)$.
The non-linear transformation is important because it affects the validity of the exclusion assumptions. Specifically, suppose that we believed that the exclusion assumptions held for a given covariate, $X_{k}$, mechanism, $M_{1}$, and outcome YY. [^14] Note that there is no guarantee there exists an outcome for which Assumptions 1-2 hold in a given system of mechanisms. Given our definition of the non-linear function $h(\cdot)$, $ADE^{\widetilde{Y}}(z,z^{\prime};X_{k}=x)=\mathbb{E}[\delta(Y(z^{\prime},M_{1},M_{2};X_{k}=x);Y(z,M_{1},M_{2};X_{k}=x)-Y(z^{\prime},M_{1},M_{2};X_{k}=x))]$. However, in general, $ADE^{\widetilde{Y}}(z,z^{\prime};X_{k}=x)\neq ADE^{\widetilde{Y}}(z,z^{\prime};X_{k}=x^{\prime})$ without other specific assumptions about the functional form of $h(\cdot)$. This means that even if Assumption 1 holds for $Y$, it generally will not hold for the transformed outcome $\widetilde{Y}$. A similar result holds for Assumption 2 and $AIE^{\widetilde{Y}}_{2}$. The logic for these observations is that the non-linear transformation “breaks” the additive separability of the mechanism from (1) other mechanism(s) and (2) predictors of the outcome. This mechanically generates additional causal moderation with different mechanisms without changing the underlying processes through which treatment affects the outcome. Given our distinction between substantive mechanisms and measurement of the effects they produce, we view the introduction of heterogeneity via a change in outcome variables as distinct from the presence of heterogeneity that exists in how mechanisms present in the world.
When are transformed outcomes used in applied work? Three cases are quite common (but non-exhaustive). The first is akin to our running example: a treatment induces a change in information or utility that then affects some discrete choice of strategy. The second holds that a treatment changes an actor’s attitude ($Y)$. But since attitudes are latent, survey researchers measure changes in the attitude by employing a Likert scale of the form:
$h(Y)$ $=\begin{cases}1&Y\in(-\infty,c_{1}]\\ 2&Y\in(c_{1},c_{2}]\\ \vdots&\\ Q&Y\in(c_{Q-1},\infty),\end{cases}$ (13)
in which $c_{t}$ denotes increasing thresholds in a latent attitude. Third, a researcher may employ non-linear transformations of an outcome to demonstrate the robustness of results to measurement choices. Here, they may bin a count variable to capture the extensive margin of some behavior or winsorize or logarithmize a skewed outcome etc. In any of these cases, even if Assumptions 1-2 were to hold for the orignal outcome (utility, attitudes, or the raw variable), they will not hold for the transformed outcome.
2.4.2 Moderators and mechanisms
Second, consider the role of a measured covariate, $X_{k}$, which we seek to use to detect the activation of focal mechanism $M_{1}$. A mechanism detector variable for $M_{1}$ is a covariate that produces different conditional AIEs at different values. To economize notation, we will denote $AIE_{j}^{Y}(X_{k}=x)=AIE_{j}^{Y}(z,z^{\prime};X_{k}=x,X_{\neg k})$ as the average indirect effect of mechanism (mediator) when $X_{k}=x$.
Definition 4.
A pre-treatment covariate $X_{k}$ is a mechanism detector variable for mechanism $j$ with respect to outcome $Y$ if for some $x,x^{\prime}\in X_{k}$, $AIE^{Y}_{j}(X_{k}=x)\neq AIE^{Y}_{j}(X_{k}=x^{\prime})$.
We then denote $\textbf{X}^{MDV}\subseteq\{X_{1},X_{2},...,X_{L}\},L\leq K$, as the (possibly empty) set of covariates that satisfy Definition 4 for the mechanism $j$. Intuitively, if $X_{k}\in\textbf{X}^{MDV}$, then covariate $X_{k}$ can serve as an indicator for a mechanism/mediator of interest. [^15] A slightly stronger version of Definition 4 holds when $Y$ is continuously differentiable with respect to $M$, $Z$, and $X_{k}$. In this case, Definition 4 can be expressed as $\frac{\partial}{\partial X_{k}}\left(\frac{\partial Y}{\partial M}\frac{\partial M}{\partial Z}\right)\neq 0$ for some $x\in X_{k}$. Under our definition of MDVs, it could be the case that $X_{k}$ moderates the effect of the treatment on the mediator. Interestingly, it could also be the case that $X_{k}$ moderates the effect of the mediator on the outcome. Both possibilities are depicted in Figure 2. Researchers using HTEs to investigate mechanisms seek to detect evidence of a mechanism using treatment-by-covariate interactions with a proposed MDV. In other words, given a proposed MDV $X_{k}$, we want to learn whether $X_{k}\in\textbf{X}^{MDV}$.
$X_{k}ZZMMYY$$X_{k}$
3 Results
We consider the conditions under which HTEs (or lack thereof) are informative about the activation of a mechanism, $M_{1}$, for some unit(s) in a sample. To do so, we analyze four exhaustive and mutually exclusive cases that vary the (1) existance of HTEs and (2) whether an outcome is non-linearly transformed or not. In each case, we will assume that Assumptions 1-2 hold for the outcome $Y$. This follows the identification result in Proposition 1. If we were not to invoke these assumptions, we could not link differences in CATEs (HTEs) to differences in the AIEs of a mechanism.
3.1 Case 1: HTEs exist for outcome $Y(Z)$
Proposition 2 analyzes the case in which HTEs exist (are observed) and Assumptions 1 and 2 are assumed to hold for the outcome, $Y$. It shows that HTEs can provide evidence that a covariate is an MDV for some mechanism of interest, $M_{1}$. Recall that if $X_{k}$ is a MDV for mechanism 1, $AIE^{Y}_{1}(X_{k}=x)\neq AIE^{Y}_{1}(X_{k}=x^{\prime})$ for some $x,x^{\prime}\in X_{k}$. This is sufficient to provide evidence that mechanism $M$ is active for at least one unit.
Proposition 2.
Suppose Assumptions 1-2 hold with respect to $X_{k}$ for outcome $Y$. If HTEs exist with respect to $X_{k}$, then $X_{k}\in\textbf{X}^{MDV}$ for mechanism $M_{1}$.
This conforms to standard interpretations that the HTEs provide evidence that a mechanism is active. Nevertheless, this finding relies critically upon the validity of exclusion assumptions, following Proposition 1. If one were to detect heterogeneity in an empirical study in this class, the heterogeneity would indeed provide evidence that the postulated mechanism, $M_{1}$, is active for some units.
3.2 Case 2: HTEs do not exist for outcome $Y(Z)$
We now consider the converse: the case when there exist no HTEs with respect to $X_{k}$ and Assumptions 1 and 2 are assumed to hold for the outcome, $Y$.
Proposition 3.
Suppose Assumptions 1-2 hold with respect to $X_{k}$ for outcome $Y$. If no HTEs exist with respect to $X_{k}$, at least one of the following must be true:
@@ 3.2 Case 2: HTEs do not exist for outcome $Y(Z)$
    $X_{k}\notin\textbf{X}^{MDV}$ for mechanism $M_{1}$.